OpenStudy (momo):

Find the area of a sector defined by a central angle of 120 degrees and a radius of 6 cm.

OpenStudy (anonymous):

\[A = (\theta/360)\Pi r^{2}\]\[A = (120/360)\Pi 6^{2}\]\[A = (1/3)\Pi36\]\[A = 12\Pi cm ^{2}\]

OpenStudy (momo):

wow that's amazing! thanks!

OpenStudy (anonymous):

no problem :)